OpenStudy (anonymous):

find three rational numbers between .1 and 3.2 ?

Parth (parthkohli):

Do you mean 0.1 and 3.2?

OpenStudy (anonymous):

oh sorry 3.1 and 3.2

Parth (parthkohli):

Oh and do they have to be exactly be in between or any rational numbers between them would work?

OpenStudy (anonymous):

yes

Parth (parthkohli):

3.15 is in between 3.1 and 3.2 Between 3.15 and 3.1 falls 3.125 Between 3.2 and 3.15 falls 3.175 so the numbers between 3.1 and 3.2 are 3.125,3.15 and 3.175

OpenStudy (anonymous):

Since you're looking for rational numbers, it may help to write your boundaries as rational numbers also. Find rational numbers between: \(\frac{31}{10} \text{ and } \frac{32}{10}\)

OpenStudy (anonymous):

n u can also tell irrational

Parth (parthkohli):

There are infinitely many rational numbers between 3.1 and 3.2, but if you mean by 3 arithmetic means, then I have given my answer.

Parth (parthkohli):

Maybe you can use 3.1223743734735725754754757435735734574575........ unending ;) and something like that.

OpenStudy (anonymous):

ok thanxx for your help

OpenStudy (anonymous):

There's a popular, and well-known irrational number between 3.1 and 3.2, commonly represented by a Greek letter ;)

OpenStudy (anonymous):

what

Parth (parthkohli):

pi!!

OpenStudy (anonymous):

it means

OpenStudy (anonymous):

see i m very poor in maths

OpenStudy (anonymous):

oh k thnxxxxxxxxx a lot

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \pi = 3.14159.... }\)

OpenStudy (anonymous):

\[\pi\], as in 3.1415926535...

OpenStudy (anonymous):

Alternatively I rather like the sequential, yet still irrational numbers of the form: \[3.101001000100001000001000000100000001000000001000000000...\]

OpenStudy (anonymous):

how would we solve it

OpenStudy (anonymous):

They are possible solutions to your questions, they do not need any extra solving :)

OpenStudy (anonymous):

ohh thanxx n bye 2 aall

OpenStudy (anonymous):

2,3 are the rational numbers